OpenStudy (anonymous):

find equation whose roots are 1/a^2 and 1/b^2 given that a & b are roots of ax^2 +bx +c

OpenStudy (anonymous):

roots mean that \[x = \frac 1{a^2}\] and \[x = \frac 1{b^2}\] that means your expression is \[(x - \frac 1{a^2})(x - \frac 1{b^2})\] expand to get your polynomial